OpenStudy (zahrah):

How many atoms of silver are in the bracelet If the silver's mass is 0.271kg?

OpenStudy (anonymous):

0.271kg = = 271g

OpenStudy (anonymous):

moles = given mass / molecular mass = 271/107

OpenStudy (anonymous):

no of atoms = n * 6.022 * 10^23 where n is no of moles

OpenStudy (zahrah):

i get it wrong

OpenStudy (zahrah):

anyone can help?

OpenStudy (anonymous):

He's saying first convert the Kg in to grams, then divide the grams of silver by it's molar mass to find the amount of moles, then using avogadros number you can convert to atoms.\[.271kg\times{1000g\over1kg}=271g\]then using\[{moles}={grams\over molar\space mass} \] We know both the molar mass and grams so solve for moles. Then after finding moles put it in the equation for finding atoms.\[moles\ \times \frac{6.022\times10^{23}atoms}{1moles}\]

OpenStudy (zahrah):

thanks